OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin 7 pi/8

OpenStudy (jdoe0001):

\(\bf \cfrac{7}{8} \times 2 \implies \cfrac{7}{4}\\ sin\left(\frac{7 \pi}{8}\right) \implies sin\left(\cfrac{\frac{7 \pi}{4}}{2}\right)\)

OpenStudy (jdoe0001):

and surely \(\bf \cfrac{7 \pi}{4}\) will be in your Unit Circle

OpenStudy (anonymous):

would it be \[1/2\sqrt{2-\sqrt{2}}\]

OpenStudy (anonymous):

as the final result

OpenStudy (jdoe0001):

\(\bf sin\left(\cfrac{\frac{7 \pi}{4}}{2}\right) \implies \sqrt{\cfrac{1-\left(-\frac{\sqrt{2}}{2}\right)}{2}}\\ \sqrt{\cfrac{1+\frac{\sqrt{2}}{2}}{2}} \implies \sqrt{\cfrac{\frac{2+\sqrt{2}}{2}}{2}} \implies \cfrac{\sqrt{2+\sqrt{2}}}{2}\)

OpenStudy (anonymous):

Thats not an answer choices